An ancient beam of wood is discovered during an archiological dig and carbon dated. They find that the sample contains 12.5% of the count rate of a similar mass of living wood. The half-life of the carbon isotope is 5600 years. Calculate the age of the ancient wood.


1
Expert's answer
2021-03-24T19:41:51-0400

Let’s use the formula for radioactive decay:


"N=N_0e^{-\lambda t},""\dfrac{N}{N_0}=e^{-\lambda t},""ln(\dfrac{N}{N_0})=ln(e^{-\lambda t}),""ln(\dfrac{N}{N_0})=-\dfrac{0.693}{T_{1/2}}t,""t=\dfrac{ln(\dfrac{N}{N_0})}{(-0.693)}T_{1/2},""t=\dfrac{ln(\dfrac{12.5\%}{100\%})}{(-0.693)}\cdot5600\ y=16803\ y."